### v1

@@ Results
We conducted experiments at the SLAC MeV-UED facility to demonstrate the feasibility of applying MOBO for online exploration and optimizations of electron beam properties in a multi-dimensional parameter space. The experimental setup and MOBO workflow are illustrated in Fig. 1. The MeV-UED beamline generates 4.2 MeV energy electron pulses using an S-band photoinjector triggered by 60 fs FWHM laser pulses. The electron pulse length downstream is sensitive to the launch phase of the electrons relative to the photoinjector RF field(referred to as the ”gun phase” and characterized by the parameter $\phi$). The electron pulses are then focused using two magnetic solenoids. One is referred to as the ”gun solenoid” and characterized by the parameter $B_{1}$, the other is referred to as the ”micro-focus solenoid” and characterized by the parameter $B_{2}$). The electrons are delivered to the Interaction Point(IP) for performing pump-probe electron diffraction experiments. The electron diffraction patterns are then recorded and used to extract fs time scale structural dynamics signals. Here, three electron beam properties are optimized. The electron pulse length($\sigma_{t}$) at IP sets the smallest temporal resolution that can be achieved for resolving fast dynamical processes. The spot size($\sigma_{x}$) at IP sets the transverse probing area and the q-resolution ($\sigma_{q}$) on the diffraction detector represents the transverse momentum resolution in the reciprocal space. Due to space charge forces, $\sigma_{t}$, $\sigma_{x}$ and $\sigma_{q}$ are competing properties that can not be simultaneously minimized. We used MOBO to explored the response of $\sigma_{t}$, $\sigma_{x}$ and $\sigma_{q}$ to the input beamline parameters [$\phi$, $B_{1}$, $B_{2}$] and obtained the PFs giving trade-offs between them. In each iteration, MOBO evaluates a candidate point in the parameter space and updates the GP surrogate model with the newly acquired data. The Expected Hypervolume Improvement (EHVI) [29] acquisition function is then used to calculate the optimal next observation point (see methods). The process is performed iteratively and ultimately providing an approximation of the best achievable PF.
@@
We first conducted $\sigma_{t}$ vs $\sigma_{q}$ optimizations by varying gun phase ($\phi$) and solenoid strengths ($B_{1}$ and $B_{2}$). The initial pulse charge was set to be 10 fC. The gun amplitude was 90 MV/m. A 200 $\mu$m collimator was used as an emittance filter located after the gun solenoid. We initialized the exploration with 10 random sampling points in the chosen subdomain of input parameter space (Table.  1). Fig. 2 illustrates the exploration result after 121 measurements. For each measurement, three observations of $\sigma_{t}$ and $\sigma_{q}$ were recorded for averaging the fluctuations due to system jitters. These measurements are then used to train GP models of $\sigma_{t}$ and $\sigma_{q}$ responses to the input parameter $\phi$, $B_{1}$ and $B_{2}$. A Matérn kernel with $\nu$ = 5/2 was used and hyperparameter training was done by maximizing the log marginal likelihood of the GP model with the gradient based Adam optimization algorithm. Measurements of $\sigma_{t}$ projected to the $\phi$ $B_{2}$ subspace, along with the posterior mean predicted by the GP model, are plotted in Fig. 2(a). Estimated length scales from the GP models are shown in Table.  1. The $\mathcal{L}$($\sigma_{t}$) for $\phi$, $B_{1}$ and $B_{2}$ are at the same scale, implying that $\sigma_{t}$ depends on all three input parameters. This is consistent with previous observations as $\sigma_{t}$ depends on both the temporal-energy chirp at gun exit(changed by $\phi$) and SC induced temporal broadening during beam transport(changed by $B_{1}$ and $B_{2}$). Fig. 2(b) and (c) illustrates the $\sigma_{t}$ and $\sigma_{q}$ responses projected to the $B_{1}$ $B_{2}$ subspace. The $\mathcal{L}$($\sigma_{q}$) for $B_{1}$ is significantly shorter than that for $B_{2}$, indicating that $\sigma_{q}$ varies faster with $B_{1}$. Again, this is consistent with previous observations that q-resolution mainly depends on the transmitted beam emittance and gun solenoid strength. On the other hand, The $\mathcal{L}$($\sigma_{q}$) for $\phi$ is significantly larger than the $\mathcal{L}$($\sigma_{q}$) for $B_{1}$ and $B_{2}$, indicating that the q-resolution is weakly correlated to gun phase. The GP predicted PF curve(denoted with green squares) is plotted in Fig. 2(b) and (c). The GP predicted PF curve connects the GP predicted optimal set points for lowest $\sigma_{t}$ at [$B_{1}$ = 0.131, $B_{2}$ = 0.240] and lowest $\sigma_{q}$ at [$B_{1}$ = 0.123, $B_{2}$ = 0.217]. Here, we define the empirical non-dominated data points as the approximate PF. The approximate PF points are plotted with green stars, the difference between approximate PF and GP predicted PF indicates that further exploration is needed to obtain the ultimate PF, and the lighter colored regions (both in $\sigma_{t}$ and $\sigma_{q}$ plots) are particularly valuable and remains to be explored. Figure. 2(d) and (e) show the calculated EHVI acquisition function projected onto the $\phi$ $B_{2}$ and $B_{1}$ $B_{2}$ space, the most valuable next observation point was found to be at [$\phi$ = 46.0, $B_{1}$ = 0.125, $B_{2}$ = 0.237]. This is consistent with the expectation that the EHVI should suggest exploring points on the predicted PF curve to efficiently improve the hypervolume. Here, the advantage of the MOBO scheme can be clearly seen. It strategically proposes the next observation point which optimally increase the Pareto Frontier HV, and is thus more data efficient than a broad, undirected search for the PF. Fig. 2(f) shows the measured data points in the objective space. The approximate PF shows clear trade-off between $\sigma_{t}$ and $\sigma_{q}$. The smallest $\sigma_{q}$ was found to be 0.162 Å-1 at $\sigma_{t}$ = 108 fs, while smallest $\sigma_{t}$ = 67 fs can be achieved with $\sigma_{q}$ = 0.217 Å-1.
@@
We then used MOBO to conduct spot size($\sigma_{x}$) vs q-resolution($\sigma_{q}$) explorations by varying $B_{1}$ and $B_{2}$ at 10 fC, 50 fC and 100 fC initial pulse charges, the input parameter ranges used are shown in Supplementary Table 1. For each run, 10 random sampling points were used to initialize the exploration and in total 150 measurements were taken. Fig. 3(a) illustrates the exploration results, all measured data points are plotted in the $\sigma_{x}$ $\sigma_{q}$ space. The approximate PFs clearly shows trade-offs between $\sigma_{x}$ and $\sigma_{q}$ at different initial pulse charges. We observe that the approximate PF obtained at lower pulse charge (10 fC) shows significant improvement compared with higher pulse charge cases (50 fC and 100 fC). This observation is consistent with previous observations that smaller $\sigma_{x}$ or $\sigma_{q}$ can be achieved by lowering initial pulse charge, as spot size and q-resolution are limited by SC induced emittance growth at 4.2 MeV beam energy. At 10 fC initial pulse charge, the smallest $\sigma_{x}$ was found to be 45 $\mu$m at $\sigma_{q}$ = 0.351 Å-1, while smallest $\sigma_{q}$ = 0.162 Å-1 can be achieved with $\sigma_{x}$ = 100 $\mu$m. We also observe that $\sigma_{q}$ flattens with $\sigma_{x}$ higher than 100 $\mu$m, implying that other source of broadening on the diffraction detector, such as the detector Point Spread Function(PSF) are the primary limitation for further improvement in the q-resolution. We ran 9 independent optimizations for the 10 fC initial pulse charge case, each with a different set of 10 randomly sampled initial points. After each measurement, the hypervolume of the approximate PF point set was calculated in normalized objective space. Figure. 3(b) shows the convergence plot of approximate PF hypervolume versus number of iterations. One can clearly see improvement of hypervolume across the optimization process. The averaged hypervolume (blue) reaches 95% of its maximum within 30 measurements. We conducted a simulated Grid Search (GS) with step size equal to 10% of the parameter space subdomain using the interpolated dataset. The hypervolume obtained using GS was 62% of that obtained using MOBO after 30 measurements and 91% after 100 measurements. A direct comparison between MOBO and GS scheme is plotted in Fig. 3(b), the advantage of MOBO to improve both search efficiency and maximum achievable hypervolume can be clearly seen.
@@ Methods
Nonparametric GP surrogate models were used to predict the value of an objective function $f(\textbf{x})$ using Bayesian statistics [27]. The GP model model includes a distribution of possible functions $y(\mathbf{x})\sim\mathcal{GP}(\mu(\mathbf{x}),k(\mathbf{x},\mathbf{x}^{\prime}))$. With $\mu(\textbf{x})$ presenting the predicted mean of objective function $y(\textbf{x})$, $k(\textbf{x},\textbf{x}^{\prime})$ is the kernel function based on the objective function behavior [20]. Given N previous measurements $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{N},y_{N})\}$ , the predictive probability distribution of the function value is given by a normal distribution:
@@
The data presented in the paper and other findings of the study are available from the corresponding author upon request.
@@ Acknowledgments
F. J., A. E. and R. J. E. designed the experiment. A. E., R. R., F. J. and C. M. wrote the software required to perform data collection and online multi-objective bayesian optimizations. F. J., A. E., X. S., S. M., S. W., D. L., M. M., A. R., M. A. K. O., P. K. and E. N. conducted the experiment. R. J. E., and M. P. M. provided guidance and oversights. F. J., R. J. E., X. S., and R. R. performed data analysis and wrote the manuscript. All authors contributed to the completion of the manuscript.
@@
Figure 2: Temporal length vs q-resolution optimization results. (a), (b) Overview of temporal length response projected to the $\phi$ $B_{2}$ and $B_{1}$ $B_{2}$ space. Gaussian Process surrogate model was constructed using measured data points(gray dots), color map(blue) represents the predicted mean of $\sigma_{t}$. (c) Overview of q-resolution response projected to the $B_{1}$ $B_{2}$ space. The color map represents the predicted mean of $\sigma_{q}$. (d), (e) On-the-fly acquisition function projected to the $\phi$ $B_{2}$ and $B_{1}$-$B_{2}$ space, colormap shows the optimal next observation regions predicted by the acquisition function. (f) Measured data points in the objective($\sigma_{t}$ $\sigma_{q}$) space, approximate Pareto Front points are highlighted with green stars.
Figure 3: Beam spot size vs q-resolution optimizations (a) Optimization results with 10 fC, 50 fC and 100 fC initial pulse charges. Measured data points are plotted in the $\sigma_{x}$-$\sigma_{q}$ space, approximate Pareto Front points are highlighted. (b) Convergence plot with averaging Pareto Front hypervolume (blue) for 9 independent optimizations (gray) for the 10 fC initial pulse charge case, shading area denotes 1 sigma variance. Interpolated grid search result (red) is plotted for comparison.

### v2

@@ Results
We conducted experiments at the SLAC MeV-UED facility to demonstrate the feasibility of applying MOBO for online exploration and optimizations of electron beam properties in a multi-dimensional parameter space. The experimental setup and MOBO workflow are illustrated in Fig. 1. The MeV-UED beamline generates 4.2 MeV energy electron pulses using an S-band photoinjector triggered by 60 fs FWHM laser pulses. The electron pulse length downstream is sensitive to the launch phase of the electrons relative to the photoinjector RF field(referred to as the ”gun phase” and characterized by the parameter $\phi$). The electron pulses are then focused using two magnetic solenoids. One is referred to as the ”gun solenoid” and characterized by the parameter $B_{1}$, the other is referred to as the ”micro-focus solenoid” and characterized by the parameter $B_{2}$. The electrons are delivered to the Interaction Point(IP) for performing pump-probe electron diffraction experiments. The electron diffraction patterns are then recorded and used to extract fs time scale structural dynamics signals. Here, three electron beam properties are optimized. The electron pulse length($\sigma_{t}$) at IP sets the smallest temporal resolution that can be achieved for resolving fast dynamical processes. The spot size($\sigma_{x}$) at IP sets the transverse probing area and the q-resolution ($\sigma_{q}$) on the diffraction detector represents the transverse momentum resolution in the reciprocal space. Due to space charge forces, $\sigma_{t}$, $\sigma_{x}$ and $\sigma_{q}$ are competing properties that can not be simultaneously minimized. We used MOBO to explored the response of $\sigma_{t}$, $\sigma_{x}$ and $\sigma_{q}$ to the input beamline parameters [$\phi$, $B_{1}$, $B_{2}$] and obtained the PFs giving trade-offs between them. In each iteration, MOBO evaluates a candidate point in the parameter space and updates the GP surrogate model with the newly acquired data. The Expected Hypervolume Improvement (EHVI) [29] acquisition function is then used to calculate the optimal next observation point (see methods). The process is performed iteratively and ultimately providing an approximation of the best achievable PF.
@@
We first conducted $\sigma_{t}$ vs $\sigma_{q}$ optimizations by varying gun phase ($\phi$) and solenoid strengths ($B_{1}$ and $B_{2}$). The initial pulse charge was set to be 10 fC. The gun amplitude was 90 MV/m. A 200 $\mu$m collimator was used as an emittance filter located after the gun solenoid. We initialized the exploration with 10 random sampling points in the chosen subdomain of input parameter space (Table  1). Fig. 2 illustrates the exploration result after 121 measurements. For each measurement, three observations of $\sigma_{t}$ and $\sigma_{q}$ were recorded for averaging the fluctuations due to system jitters. These measurements are then used to train GP models of $\sigma_{t}$ and $\sigma_{q}$ responses to the input parameter $\phi$, $B_{1}$ and $B_{2}$. A Matérn kernel with $\nu$ = 5/2 was used and hyperparameter training was done by maximizing the log marginal likelihood of the GP model with the gradient based Adam optimization algorithm. Measurements of $\sigma_{t}$ projected to the $\phi$ $B_{2}$ subspace, along with the posterior mean predicted by the GP model, are plotted in Fig. 2(a). Estimated length scales from the GP models are shown in Table  1. The $\mathcal{L}$($\sigma_{t}$) for $\phi$, $B_{1}$ and $B_{2}$ are at the same scale, implying that $\sigma_{t}$ depends on all three input parameters. This is consistent with previous observations as $\sigma_{t}$ depends on both the temporal-energy chirp at gun exit(changed by $\phi$) and SC induced temporal broadening during beam transport(changed by $B_{1}$ and $B_{2}$). Fig. 2(b) and (c) illustrates the $\sigma_{t}$ and $\sigma_{q}$ responses projected to the $B_{1}$ $B_{2}$ subspace. The $\mathcal{L}$($\sigma_{q}$) for $B_{1}$ is significantly shorter than that for $B_{2}$, indicating that $\sigma_{q}$ varies faster with $B_{1}$. Again, this is consistent with previous observations that q-resolution mainly depends on the transmitted beam emittance and gun solenoid strength. On the other hand, The $\mathcal{L}$($\sigma_{q}$) for $\phi$ is significantly larger than the $\mathcal{L}$($\sigma_{q}$) for $B_{1}$ and $B_{2}$, indicating that the q-resolution is weakly correlated to gun phase. The GP predicted PF curve(denoted with green squares) is plotted in Fig. 2(b) and (c). The GP predicted PF curve connects the GP predicted optimal set points for lowest $\sigma_{t}$ at [$B_{1}$ = 0.131, $B_{2}$ = 0.240] and lowest $\sigma_{q}$ at [$B_{1}$ = 0.123, $B_{2}$ = 0.217]. Here, we define the empirical non-dominated data points as the approximate PF. The approximate PF points are plotted with green stars, the difference between approximate PF and GP predicted PF indicates that further exploration is needed to obtain the ultimate PF, and the lighter colored regions (both in $\sigma_{t}$ and $\sigma_{q}$ plots) are particularly valuable and remains to be explored. Fig. 2(d) and (e) show the calculated EHVI acquisition function projected onto the $\phi$ $B_{2}$ and $B_{1}$ $B_{2}$ space, the most valuable next observation point was found to be at [$\phi$ = 46.0, $B_{1}$ = 0.125, $B_{2}$ = 0.237]. This is consistent with the expectation that the EHVI should suggest exploring points on the predicted PF curve to efficiently improve the hypervolume. Here, the advantage of the MOBO scheme can be clearly seen. It strategically proposes the next observation point which optimally increase the Pareto Frontier HV, and is thus more data efficient than a broad, undirected search for the PF. Fig. 2(f) shows the measured data points in the objective space. The approximate PF shows clear trade-off between $\sigma_{t}$ and $\sigma_{q}$. The smallest $\sigma_{q}$ was found to be 0.162 Å-1 at $\sigma_{t}$ = 108 fs, while smallest $\sigma_{t}$ = 67 fs can be achieved with $\sigma_{q}$ = 0.217 Å-1.
@@
We then used MOBO to conduct spot size($\sigma_{x}$) vs q-resolution($\sigma_{q}$) explorations by varying $B_{1}$ and $B_{2}$ at 10 fC, 50 fC and 100 fC initial pulse charges, the input parameter ranges used are shown in Supplementary Table 1. For each run, 10 random sampling points were used to initialize the exploration and in total 150 measurements were taken. Fig. 3(a) illustrates the exploration results, all measured data points are plotted in the $\sigma_{x}$ $\sigma_{q}$ space. The approximate PFs clearly shows trade-offs between $\sigma_{x}$ and $\sigma_{q}$ at different initial pulse charges. We observe that the approximate PF obtained at lower pulse charge (10 fC) shows significant improvement compared with higher pulse charge cases (50 fC and 100 fC). This observation is consistent with previous observations that smaller $\sigma_{x}$ or $\sigma_{q}$ can be achieved by lowering initial pulse charge, as spot size and q-resolution are limited by SC induced emittance growth at 4.2 MeV beam energy. At 10 fC initial pulse charge, the smallest $\sigma_{x}$ was found to be 45 $\mu$m at $\sigma_{q}$ = 0.351 Å-1, while smallest $\sigma_{q}$ = 0.162 Å-1 can be achieved with $\sigma_{x}$ = 100 $\mu$m. We also observe that $\sigma_{q}$ flattens with $\sigma_{x}$ higher than 100 $\mu$m, implying that other source of broadening on the diffraction detector, such as the detector Point Spread Function(PSF) are the primary limitation for further improvement in the q-resolution. We ran 9 independent optimizations for the 10 fC initial pulse charge case, each with a different set of 10 randomly sampled initial points. After each measurement, the hypervolume of the approximate PF point set was calculated in normalized objective space. Fig. 3(b) shows the convergence plot of approximate PF hypervolume versus number of iterations. One can clearly see improvement of hypervolume across the optimization process. The averaged hypervolume (blue) reaches 95% of its maximum within 30 measurements. We conducted a simulated Grid Search (GS) with step size equal to 10% of the parameter space subdomain using the interpolated dataset. The hypervolume obtained using GS was 62% of that obtained using MOBO after 30 measurements and 91% after 100 measurements. A direct comparison between MOBO and GS scheme is plotted in Fig. 3(b), the advantage of MOBO to improve both search efficiency and maximum achievable hypervolume can be clearly seen.
@@ Methods
Nonparametric GP surrogate models were used to predict the value of an objective function $f(\textbf{x})$ using Bayesian statistics [27]. The GP model includes a distribution of possible functions $y(\mathbf{x})\sim\mathcal{GP}(\mu(\mathbf{x}),k(\mathbf{x},\mathbf{x}^{\prime}))$. With $\mu(\textbf{x})$ presenting the predicted mean of objective function $y(\textbf{x})$, $k(\textbf{x},\textbf{x}^{\prime})$ is the kernel function based on the objective function behavior [20]. Given N previous measurements $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{N},y_{N})\}$ , the predictive probability distribution of the function value is given by a normal distribution:
@@
The raw data generated in this study have been deposited in the Zenodo repository under accession code: https://doi.org/10.5281/zenodo.11095450[44]
@@ Acknowledgments
F. J., A. E. and R. J. E. designed the experiment. A. E., R. R., F. J. and C. M. wrote the software required to perform data collection and online multi-objective bayesian optimizations. F. J., A. E., X. S., S. M., S. W., D. L., M. M., A. R., M. A. K. O., P. K., X. W. and E. N. conducted the experiment. R. J. E. and M. M. provided guidance and oversights. F. J., R. J. E., X. S. and R. R. performed data analysis and wrote the manuscript. All authors contributed to the completion of the manuscript.
@@
Figure 2: Temporal length vs q-resolution optimization results. (a), (b) Overview of temporal length response projected to the $\phi$ $B_{2}$ and $B_{1}$ $B_{2}$ space. Gaussian Process surrogate model was constructed using measured data points(gray dots), color map(blue) represents the predicted mean of $\sigma_{t}$. (c) Overview of q-resolution response projected to the $B_{1}$ $B_{2}$ space. The color map represents the predicted mean of $\sigma_{q}$. (d), (e) On-the-fly acquisition function projected to the $\phi$ $B_{2}$ and $B_{1}$-$B_{2}$ space, colormap shows the optimal next observation regions predicted by the acquisition function. (f) Measured data points in the objective($\sigma_{t}$ $\sigma_{q}$) space, approximate Pareto Front points are highlighted with green stars. The corresponding hypervolume convergence plot is shown in Supplementary Fig. 1.
Figure 3: Beam spot size vs q-resolution optimizations (a) Optimization results with 10 fC, 50 fC and 100 fC initial pulse charges. Measured data points are plotted in the $\sigma_{x}$-$\sigma_{q}$ space, approximate Pareto Front points are highlighted. (b) Convergence plot with averaging Pareto Frontier hypervolume (blue) for 9 independent optimizations (gray) for the 10 fC initial pulse charge case, shading area denotes 1 sigma variance. Interpolated grid search result (red) is plotted for comparison.